# ALMA chemical survey of disk-outflow sources in Taurus (ALMA-DOT)

### v2

@@ Abstract
Aims. The outermost component of the wind traced by CO emission is analyzed in detail to determine its relationship with the disk and its substructures.
Methods. A spectro-imaging investigation is conducted using ALMA observations of the 12CO (2-1) line at 1.3 mm, with 0.2 km sand spectral and angular resolution, in the framework of the ALMA-DOT project. Relevant wind parameters are derived, allowing a tomographic reconstruction of the morphology and kinematics of the red-shifted lobe of the outflow to be compared with theoretical models.
Results. The data channel maps and position-velocity diagrams show a rich substructure of concatenated bubble- and arc-shaped features, whose size and distance from the source continuously increase with velocity. The superposition of such features generates the apparent conical shape. The spatio-kinematical properties suggest that the flow presents distinct nested shells with increasing velocity and steeper velocity increase with distance from the star going toward the axis, and rotating in the same sense of the disk. The wind parameters are compared with the predictions of magnetohydrodynamic (MHD) disk winds. Under this hypothesis, the launch radii of the three outermost shells are found to cluster around the positions of three rings in the dust distribution of the disk located at 58, 72 and 86 au from the star. We derive a magnetic lever arm , higher than that commonly adopted in models of MHD winds from the outer disk. Interpretations are discussed.
Conclusions. The properties of the CO outflow from HL Tau appear to be compatible with magnetized disk winds with launch radii in the region at 50 to 90 au from the source. As such, the wind may be capable of removing angular momentum also from the outer disk. The arrangement of the wind in nested shells with brighter emission rooted at the location of ring substructures could support the results of recent non-ideal MHD simulations according to which magnetic instabilities can spontaneously generate both the ring-gap system and a connected inhomogeneous layered wind, alternatively to the action of yet undetected protoplanets. Further observational analyses and comparisons with other classes of models will help establish the role of magnetic effects in the process of planet formation.
@@
One of the most spectacular discoveries of the last decade has been the detection in the continuum emission at millimeter wavelengths of concentric gaps and rings in disks around young stars. The first example found is the disk around HL Tau. Observed with the Atacama Large Millimeter Array (ALMA)111https://www.eso.org/sci/facilities/alma.html, it showed in exquisite detail a sequence of concentric rings and gaps (ALMA Partnership et al., 2015), now commonly seen in many other disks (e.g., Andrews et al. 2018).
@@
The first interpretation offered was that these gaps are carved by young massive planets embedded in the disk (e.g., Lodato et al. 2019). However, this scenario raises open issues, such as the amount of mass necessary to produce massive planets at such large distances from the star (up to 100 au in HL Tau) and in such short timescales (less than 1 Myr). In addition, the predicted embedded planets inside the gaps remain elusive so far, except for PDS 70, where the ”gap” actually is a large inner cavity (Keppler et al., 2018).
@@
Several other disk processes are able to produce ring/gap structures. Perhaps the most generic are magnetic instabilities in the low-ionization ”dead zone” (e.g., Suriano et al., 2018; Riols & Lesur, 2019). This process could be widespread in young disks that harbor a residual primordial magnetic field. In this case, an inhomogeneous magnetohydrodynamic (MHD) disk wind (DW) is created, and the disk self-organizes in a system of rings and gaps, with denser wind layers launched from the rings and less dense ones from the gaps (e.g. Suriano et al. 2019; Riols et al. 2020). An asymmetry also develops between the red and blue lobes (Béthune et al., 2017), as observed in almost all cases (e.g. Podio et al. 2011, 2021). Finding firm evidence for such layered MHD disk winds is thus an essential step in a clarification of the origin of disk substructures and of the role of magnetic fields in planet formation.
@@
Signatures of layered outflows have indeed been found with ALMA in a growing number of young systems, in the form of rotating molecular winds flowing in a cavity of apparent conical shape (e.g., TMC1A, (Bjerkeli et al., 2016); HH 46/47, (Zhang et al., 2016, 2019); HH212 (Lee et al., 2017); DO Tau, (Fernández-López et al., 2020); HH 30, (Louvet et al., 2018; López-Vázquez et al., 2024; Ai et al., 2024); DG TauB (Zapata et al., 2015; de Valon et al., 2020, 2022); HH270mms1-A, (Omura et al., 2024); L1448-mm, (Nazari et al., 2024)). The CO outflow is associated in some cases with a coaxial molecular wind that emits in lines, with a fast collimated atomic jet flowing along the axis of the system (e.g., HH 46/47, (Nisini et al., 2024), DG Tau B (Delabrosse et al., 2024)). The origin of the layered structure in these CO winds is debated. In addition to the magnetic instabilities mentioned above, other studies invoke the formation of swept-up wind shells (Zhang et al., 2019), peculiarities of the MHD disk-wind accretion/ejection engine (Louvet et al., 2018; de Valon et al., 2020), flow wiggling (de Valon et al., 2022), the action of the wings of subsequent bow shocks of time-variable jets (Tabone et al., 2018; Rabenanahary et al., 2022), or the evolution of a unified wide-angle magnetized flow with a jet-bearing X-wind (Shang et al., 2020, 2023; Ai et al., 2024).
@@
Among the observable sources, HL Tau stands out as the first and only object where a possible coincidence between the wind structure and the distribution of disk ring/gaps can be studied. In fact, in the other cases the flow is associated with a disk where no gaps and rings are detected. In some cases, this can be due to the geometry of the observation (for example, the edge-on disks of HH 30). In general, however, the accretion/ejection activity associated with young stellar objects is expected to decrease with the source evolutionary stage; hence powerful and bright molecular winds and jets are associated with Class 0 protostars (e.g., Lee et al. 2018; Tabone et al. 2017; Podio et al. 2021) while only faint molecular emission is associated with evolved Class II stars. On the other hand, substructures such as rings and gaps are commonly detected in pre-main sequence disks (e.g., SHARP LP, Andrews et al. 2018), while they are rarely detected in younger disks (see, e.g., Segura-Cox et al. 2020; Sheehan & Eisner 2017, 2018; de Valon et al. 2020). In this context, HL Tau is intermediate between an evolved accreting protostar (Class I) and a classical TTauri star (Class II), and is associated with both a disk with clearly detected rings/gaps and a bright extended and structured outflow. HL Tau is therefore an ideal target for investigating the possible relationship between the disk substructures and the wind properties and for testing the proposed scenarios for their formation.
@@
The central object, a young star of mass 2.1 M⊙ (Yen et al., 2019a), is the source of a composite outflow seen at optical, infra-red, and mm wavelengths. The collimated atomic jet that shines in shock-excited lines such as [S II]6716,6731, H, and [O I]6300, stretches out of the reflection nebula surrounding the star for about 1.5 along the direction PA 50°, with an inclination angle with respect to the line of sight of (Mundt et al., 1990; Anglada et al., 2007; Pyo et al., 2006; Movsessian et al., 2012). The jet is compatible with being perpendicular to the disk, which has an inclination angle with respect to the line of sight = 46.2 (defined as for face-on), as determined in ALMA Partnership et al. (2015) from the data at 02 resolution. The same study reports that the major axis of the disk lies along PAdisk = +138.2 measured from north to east.
@@
The system was imaged at 02 resolution with adaptive optics in the near-infrared, showing a collimated jet in [FeII] 1.64 m and a surrounding wider angle warm wind emitting in H2 2.12 m , that flows in a conical cavity detected in scattered light (Takami et al., 2007; Beck et al., 2008). A coaxial molecular outflow has been observed in the CO(3) and CO(1) lines (Lumbreras & Zapata, 2014; ALMA Partnership et al., 2015; Klaassen et al., 2016). The data show a prominent conical redshifted flow to the south-west of HL Tau, with the axis almost aligned with the atomic jet, and with a full opening angle starting at 90°and then narrowing at 60°further from the source (Klaassen et al., 2016). The inclination with respect to the line of sight is assumed to be along the perpendicular to the disk (i.e., =). The north-eastern lobe of the molecular outflow is much fainter and appears as a wide cavity spread within the ambient.
@@
In this paper, we investigate the structure of the CO outflow in more detail analyzing a dataset obtained within the ALMA-DOT program (ALMA chemical survey of Disk-Outflow sources in Taurus), an observational campaign aimed at characterizing the gas in disks and outflows of embedded sources in the Taurus star-forming region. The high angular/spectral resolution of ALMA allows us to disentangle and characterize the emission from the various elements of the systems (disk, outflows, streamers, envelope). Previous papers in the ALMA-DOT series focus on the chemical properties of the disks (Podio et al., 2019, 2020a, 2020b; Codella et al., 2020) and of the associated streamers and outflows (Garufi et al., 2020, 2022), while Garufi et al. (2021) describes the collective properties of the targets examined as a whole. The present work focuses on the observations of the molecular outflow around HL Tau, as imaged in the line of 12CO (2) at 1.3 mm with a resolution of 03 and 0.2 km s, and on the implications of these findings for the structure of the disk.
@@
The paper is organized as follows. Section 2 details the observations and the reduction of the data. Section 3 describes the results obtained, while the derived hypothesis for the structure of the wind is described in Section 4. In Section 5 we discuss the comparison of the results with theoretical models in an MHD disk wind scenario. The conclusions of this study and the perspectives for future analyses are outlined in Section 6.
@@ 2 Observations and data reduction
Cycle 6 observations of the ALMA-DOT campaign (Podio et al., 2019) were taken with ALMA-Band 6 on October 28, 2018 in an extended configuration with baselines ranging from 15 m to 1.4 km (project 2018.1.01037.S, PI: L. Podio). The field of view is a circular region around HL Tau of radius 185, corresponding to about 2.7103 au at the distance from the source, 147.3 pc (Galli et al., 2018). The integration time sums up to 110 minutes (Garufi et al., 2021). The bandpass and phase calibrators were J0423-0120 and J0510+1800, respectively. The correlator setup consisted of high-resolution (0.141 MHz) spectral windows (SPWs) covering 12CO (2), CN (2), o-H2CO (3 CS (5), and CH3OH () (A). The molecular parameters were taken from the Cologne Database of Molecular Spectroscopy (CDMS) (Müller et al., 2001). In this paper, we focus on the 12CO (2) emission (rest frequency 230.538 GHz), while the description in other molecular lines can be found in Garufi et al. (2020).
Data were reduced using the standard procedure using the Common Astronomy Software Applications package (CASA, McMullin et al. 2007) version 4.7.2. Self-calibration was performed on the continuum emission and applied on the line-free continuum, and the continuum subtracted line emission. The procedure improved the continuum S/N by a factor of 3.3. The final continuum maps and spectral cubes were produced with tclean by applying a manually selected mask to the signal and iterating until the residuals revealed no significant source emission. We adopted Briggs weighting with robust=0.0 for the bright CO line, to maximize angular resolution, and set a channel width of 0.2 km s. The clean beam of the self-calibrated maps in CO is 0.26″0.31″ with a PA of -3.12°. The r.m.s. noise per channel is mJy beam-1. The beam in the continuum images is 0.23″0.26″ with PA= -6.77°. The r.m.s. for the continuum image is 0.12 mJy beam-1.
@@ 3.1 Continuum emission
The continuum at 1.3 mm has an integrated flux of 844.9 1.9 mJy, with peak intensity of 95.48 0.19 mJy beam-1 located at R.A. 04h 31m 38s.43, and Dec. 18∘ 13′ 5717. At our angular resolution the intensity distribution is smooth, and the sequence of rings and gaps is not visible. The FWHM along the major and minor axes, determined with a 2D Gaussian fit deconvolved from the beam, is and mas, respectively. From these values, we estimate an inclination of the disk (with 0° corresponding to face-on). The fit gives a disk position angle, PAdisk = +138.25. Both values agree with those reported in ALMA Partnership et al. (2015).
@@ 3.2 CO (2) emission
The ALMA datacube in the CO (2 line consists of 290 velocity channels with VLSR ranging from -26 to +32 km s, in steps of 0.2 km s. The emission of the outflow is detected from -10 km sto +31.8 km s. The systemic velocity is Vsys=+7.1 km s(Garufi et al., 2021).
@@
Figure 1 presents the intensity integrated over the velocity (moment-0 map) in the full field of view. The left panel shows the emission integrated over the blue-shifted velocity range V, while the integration in the right panel is for VLSR velocities between +8 and +30 km s, illustrating the red-shifted emission. The faint blue-shifted emission is distributed along a wide-angle cavity of semi-spherical shape pointing towards the north-east (NE), with a distinct bright feature on the northern side. The emission in the red-shifted south-west (SW) lobe presents two main components, an inner flow of apparent conical shape around an axis oriented at PA = 228°(coinciding with the minor axis of the disk) with full opening angle starting off at 90°and then narrowing at 60°further from the source (Klaassen et al., 2016) and an outer faint conical wide wind, surrounding in projection both the disk and the base of the inner outflow, with vertex at 15 NE of the source and opening angle of 120° . To guide the eye, the approximate spatial limits of these components are indicated by solid white curves in the figure.222 The opening angle was determined with a linear fit to the wind edges, identified by the position of the zeroes of the second-order derivative of the emission along cuts parallel to axis, in the first 2″of propagation. A molecular emission ridge is evident at 14” from the source. This is part of a wider elliptical feature, as shown in ALMA Partnership et al. (2015), Fig. 1, and reported here with a dashed white elliptical curve.
@@
The intensity-weighted mean velocity (moment-1 map) of the emission is presented in Fig. 2. The NE sector hosts blue-shifted material that is not collimated, while the red-shifted SW lobe is structured. The transition in velocity between the two wind components of the red lobe identified in the moment-0 map shows that the outer wide wind is slower than 10 km s, while the inner flow has increasing velocities from the border toward the axis, up to about 30 km s.
@@
The asymmetry between the blue- and red-shifted lobes may be an effect of the inclination angle with respect to the plane of the sky. For example, Klaassen et al. (2016) illustrates how the large aperture of the initial portion of the HL Tau flow causes the far side of the approaching lobe to be imaged at red-shifted velocities close to the systemic. However, while this effect may explain the loss of the upper part of the large arc at blue-shifted velocities, it would require the approaching lobe to have an opening angle larger than the receding one to explain the absence of a well-developed conical structure. A trace of faint conical arms of aperture angle similar to that in the red lobe is barely seen at the base of the blue-shifted lobe, but it is not seen further out, and it is covered by the signal of the large semicircular arc. This structure of the blue lobe may be due to HL Tau being located on the wall of a large-scale molecular bubble inflated by nearby XZ Tau, at 23″east of the source (Welch et al., 2000; Yen et al., 2019b). The expanding bubble may have compressed the outer envelope of HL Tau on its NE side, smashing the CO cavity, while the SW lobe remained unaffected. Signatures of such a large-scale bubble are visible at low blue-shifted velocities in Fig. 2.
@@
In the following we concentrate the analysis on the red-shifted lobe that presents a well-defined and relatively bright structure over a wide range of velocities. This allows a detailed inspection of the spatio-kinematic properties, for a meaningful comparison with proposed acceleration mechanisms. We defer the study of the fainter blue-shifted lobe to a forthcoming investigation, adding that the squared 10″at the center of the region were discussed in Garufi et al. (2022), where the kinematics of the rotating disk is analyzed from the same dataset, as well as possible accretion streamers.

### v3

@@ Abstract
Aims. The outermost component of the wind traced by CO emission was analyzed in detail to determine its relationship with the disk and its substructures.
Methods. A spectro-imaging investigation was conducted using ALMA observations of the 12CO (2-1) line at 1.3 mm, with 0.2 km sand spectral and angular resolution, in the framework of the ALMA-DOT project. The relevant wind parameters were derived, allowing a tomographic reconstruction of the morphology and kinematics of the redshifted lobe of the outflow to be compared with theoretical models.
Results. The data channel maps and position-velocity diagrams show a rich substructure of concatenated bubble- and arc-shaped features, whose size and distance from the source continuously increase with velocity. The superposition of such features generates the apparent conical shape. The spatial-kinematic properties suggest that the flow presents distinct nested shells with a progressively higher propagation speed and a steeper speed gradient with distance from the source for the shells closest to the axis, and rotating in the same direction as the disk. The wind parameters were compared with the predictions of magnetohydrodynamic (MHD) disk winds. Under this hypothesis, the launch radii of the three outermost shells are found to correspond to the positions of three rings in the dust emission distribution of the disk located at 58, 72, and 86 astronomical units from the star. We derive a magnetic lever arm , higher than that commonly adopted in models of MHD winds from the outer disk. Interpretations are discussed.
Conclusions. The properties of the CO outflow from HL Tau appear to be compatible with magnetized disk winds with launch radii in the region at 50 to 90 au from the source. As such, the wind may be capable of removing angular momentum also from the outer disk. The arrangement of the wind in nested shells with brighter emission rooted at the location of ring substructures could support the results of recent non-ideal MHD simulations according to which magnetic instabilities can spontaneously generate both the ring-gap system and a connected inhomogeneous layered wind, alternatively to the action of protoplanets. Further observational analyses and comparisons with other classes of models will help establish the role of magnetic effects in the process of planet formation.
@@
One of the most spectacular discoveries of the last decade has been the detection in the continuum emission at millimeter wavelengths of concentric gaps and rings in disks around young stars. The first example found was the disk around HL Tau. Observed with the Atacama Large Millimeter Array (ALMA),111https://www.eso.org/sci/facilities/alma.html it showed in exquisite detail a sequence of concentric rings and gaps (ALMA15), now commonly seen in many other disks (e.g., Andrews18).
@@
The first interpretation offered was that these gaps are carved by young massive planets embedded in the disk (e.g., Lodato19). However, this scenario raises open issues, such as the amount of mass necessary to produce massive planets at such large distances from the star (up to 100 au in HL Tau) and on such short timescales (less than 1 Myr). In addition, the predicted embedded planets inside the gaps remain elusive, except for the cases of PDS 70 (Keppler18) and WISPIT 2 (vanCapelleveen25), where the gap actually is a large inner cavity.
@@
Several other disk processes are able to produce ring–gap structures. Perhaps the most generic are magnetic instabilities in the low-ionization dead zone (e.g., Suriano18; Riols19). This process could be widespread in young disks that harbor a residual primordial magnetic field. In this case, an inhomogeneous magnetohydrodynamic (MHD) disk wind (DW) is created, and the disk self-organizes in a system of rings and gaps, with denser wind layers launched from the rings and less dense ones from the gaps (e.g., Suriano19; Riols20). An asymmetry also develops between the red and blue lobes (Bethune17), as observed in almost all cases (e.g., Podio11; Podio21). Finding firm evidence of such layered MHD disk winds is thus an essential step in a clarification of the origin of disk substructures and of the role of magnetic fields in planet formation.
@@
Signatures of layered outflows have indeed been found with ALMA in a growing number of young systems, in the form of rotating molecular winds flowing in a cavity of apparent conical shape (e.g., TMC1A, (Bjerkeli16); HH 46/47, (Zhang16; Zhang19); HH212 (Lee17); DO Tau, (Fernandez20); HH 30, (Louvet18; Lopez24; Ai24); DG Tau B (Zapata15; DeValon20; DeValon22); HH270mms1-A, (Omura24); L1448-mm, (Nazari24)). The CO outflow is associated in some cases with a coaxial molecular wind that emits in lines, with a fast collimated atomic jet flowing along the axis of the system (e.g., HH 46/47, (Nisini24), DG Tau B (Delabrosse24)). The origin of the layered structure in these CO winds is debated. In addition to the magnetic instabilities mentioned above, other studies invoke the formation of swept-up wind shells (Zhang19), peculiarities of the MHD disk-wind accretion–ejection engine (Louvet18; DeValon20), flow wiggling (DeValon22), the interaction of a pulsed jet with its environment (Tabone18), jet-driven wide bow shocks expanding in a stratified medium (Rabenanahary22), composite scenarios including entrainment of the envelope material, rotating winds and inner shocked winds (Lopez24b), or the evolution of a unified wide-angle magnetized flow with a jet-bearing X-wind (Shang20; Shang23; Ai24).
@@
Among the observable sources, HL Tau stands out as the first and only object where a possible coincidence between the wind structure and the distribution of disk rings and gaps can be studied. In the other cases the flow is associated with a disk where no gaps and rings are detected. In some cases, this can be due to the geometry of the observation (for example, the edge-on disks of HH 30). In general, however, the accretion–ejection activity associated with young stellar objects is expected to decrease with the source evolutionary stage; hence, powerful and bright molecular winds and jets are associated with Class 0 protostars (e.g., Lee18; Tabone17; Podio21), while only faint molecular emission is associated with evolved Class II stars. On the other hand, substructures such as rings and gaps are commonly detected in pre-main sequence disks (e.g., SHARP LP, Andrews18), while they are rarely detected in younger disks (see, e.g., Segura-Cox20; sheehan17; sheehan18; DeValon20). In this context, HL Tau is intermediate between an evolved accreting protostar (Class I) and a classical T Tauri star (Class II), and is associated with both a disk with clearly detected rings and gaps and a bright extended and structured outflow. HL Tau is therefore an ideal target for investigating the possible relationship between the disk substructures and the wind properties and for testing the proposed scenarios for their formation.
@@
The central object, a young star of mass 2.1 M⊙ (Yen19a), is the source of a composite outflow seen at optical, infrared, and millimeter wavelengths. The collimated atomic jet that shines in shock-excited lines such as [S II]6716,6731, H, and [O I]6300, stretches out of the reflection nebula surrounding the star for about 1.5 along the direction PA 50°, with an inclination angle with respect to the line of sight of (Mundt90; Anglada07; Pyo06; Movsessian12). The jet is compatible with being perpendicular to the disk, which has an inclination angle with respect to the line of sight = 46.2 (defined as for face-on), as determined in ALMA15 from the data at 02 resolution. The same study reports that the major axis of the disk lies along PAdisk = +138.2 measured from north to east.
@@
The system was imaged at 02 resolution with adaptive optics in the near-infrared, showing a collimated jet in [FeII] 1.64 m and a surrounding wider-angle warm wind emitting in H2 2.12 m that flows in a conical cavity detected in scattered light (Takami07; Beck08). A coaxial molecular outflow has been observed in the CO(3) and CO(1) lines (Lumbreras14; ALMA15; Klaassen16). The data show a prominent conical redshifted flow to the southwest of HL Tau, with the axis almost aligned with the atomic jet, and with a full opening angle starting at 90°and then narrowing at 60°farther from the source (Klaassen16). The inclination with respect to the line of sight is assumed to be along the perpendicular to the disk (i.e. =). The northeastern lobe of the molecular outflow is much fainter and appears as a wide cavity spread within the ambient.
@@
For this paper we investigated the structure of the CO outflow in greater detail by analyzing a dataset obtained within the ALMA chemical survey of Disk-Outflow sources in Taurus (ALMA-DOT) program, an observational campaign aimed at characterizing the gas in disks and outflows of embedded sources in the Taurus star-forming region. The high angular and spectral resolution of ALMA allows us to disentangle and characterize the emission from the various elements of the systems (disk, outflows, streamers, envelope). Previous papers in the ALMA-DOT series focus on the chemical properties of the disks (Podio19; Podio20a; Podio20b; Codella20) and of the associated streamers and outflows (Garufi20; Garufi22), while Garufi21 describes the collective properties of the targets examined as a whole. The present work focuses on the observations of the molecular outflow around HL Tau, as imaged in the line of 12CO (2) at 1.3 mm with a resolution of 03 and 0.2 km s, and on the implications of these findings for the structure of the disk.
@@
The paper is organized as follows. Section 2 details the observations and the reduction of the data. Section 3 describes the results obtained, while the derived hypothesis for the structure of the wind is described in Section 4. In Section 5 we discuss the comparison of the results with theoretical models in a MHD disk wind scenario. Our conclusions of this study and the perspectives for future analyses are outlined in Section 6.
@@ 2 Observations and data reduction
Cycle 6 observations of the ALMA-DOT campaign (Podio19) were taken with ALMA-Band 6 on October 28, 2018 in an extended configuration with baselines ranging from 15 m to 1.4 km (project 2018.1.01037.S, PI: L. Podio). The field of view was a circular region around HL Tau of radius 185, corresponding to about 2.7103 au at the distance from the source, 147.3 pc (Galli18). The integration time summed up to 110 minutes (Garufi21). The bandpass and phase calibrators were J0423-0120 and J0510+1800, respectively. The correlator setup consisted of high-resolution (0.141 MHz) spectral windows (SPWs) covering 12CO (2), CN (2), o-H2CO (3 CS (5), and CH3OH () (A). The molecular parameters were taken from the Cologne Database of Molecular Spectroscopy (CDMS) (muller01). In this paper, we focus on the 12CO (2) emission (rest frequency 230.538 GHz), while the description in other molecular lines can be found in Garufi20.
Data were reduced using the standard procedure using the Common Astronomy Software Applications package (CASA, McMullin07) version 4.7.2. Self-calibration was performed on the continuum emission and applied on the line-free continuum, and the continuum subtracted line emission. The procedure improved the continuum S/N by a factor of 3.3. The final continuum maps and spectral cubes were produced with tclean by applying a manually selected mask to the signal and iterating until the residuals revealed no significant source emission. We adopted Briggs weighting with robust=0.0 for the bright CO line, to maximize angular resolution, and set a channel width of 0.2 km s. The clean beam of the self-calibrated maps in CO is 0.26″0.31″ with a PA of -3.12°. The r.m.s. noise per channel is mJy beam-1. The beam in the continuum images is 0.23″0.26″ with PA= -6.77°. The r.m.s. for the continuum image is 0.12 mJy beam-1.
@@ 3.1 Continuum emission
The continuum at 1.3 mm has an integrated flux of 844.9 1.9 mJy, with peak intensity of 95.48 0.19 mJy beam-1 located at R.A. 04h 31m 38s.43, and Dec. 18∘ 13′ 5717. At our angular resolution the intensity distribution is smooth, and the sequence of rings and gaps is not visible. The FWHM along the major and minor axes, determined with a 2D Gaussian fit deconvolved from the beam, is and mas, respectively. From these values, we estimated an inclination of the disk (with 0° corresponding to face-on). The fit gives a disk position angle, PAdisk = +138.25. Both values agree with those reported in ALMA15.
@@ 3.2 CO (2) emission
The ALMA datacube in the CO (2 line consists of 290 velocity channels with VLSR ranging from -26 to +32 km s, in steps of 0.2 km s. The emission of the outflow is detected from -10 km sto +31.8 km s. The systemic velocity is Vsys=+7.1 km s(Garufi21).
@@
Figure 1 presents the intensity integrated over the velocity (moment-0 map) in the full field of view. The left panel shows the emission integrated over the blueshifted velocity range V, while the integration in the right panel is for VLSR velocities between +8 and +30 km s, illustrating the redshifted emission. The faint blueshifted emission is distributed along a wide-angle cavity of semi-spherical shape pointing toward the northeast (NE), with a distinct bright feature on the northern side. The emission in the redshifted southwest (SW) lobe presents two main components, an inner flow of apparent conical shape around an axis oriented at PA = 228°(coinciding with the minor axis of the disk) with a full opening angle starting at 90°and then narrowing at 60°farther from the source (Klaassen16) and an outer faint conical wide wind, in projection surrounding both the disk and the base of the inner outflow, with vertex at 15 NE of the source and opening angle of 120° . To guide the eye, the approximate spatial limits of these components are indicated by solid white curves in the figure.222The opening angle was determined with a linear fit to the wind edges, identified by the position of the zeroes of the second-order derivative of the emission along cuts parallel to axis, in the first 2″of propagation.
A molecular emission ridge is evident at 14” from the source. This is part of a wider elliptical feature, as shown in ALMA15, Fig. 1, and reported here with a dashed white elliptical curve.
@@
The intensity-weighted mean velocity (moment-1 map) of the emission is presented in Fig. 2. The NE sector hosts blueshifted material that is not collimated, while the redshifted SW lobe is structured. The transition in velocity between the two wind components of the red lobe identified in the moment-0 map shows that the outer wide wind is slower than 10 km s, while the inner flow has increasing velocities from the border toward the axis, up to about 30 km s.
@@
The asymmetry between the blue- and redshifted lobes may be an effect of the inclination angle with respect to the plane of the sky. For example, Klaassen16 illustrates how the large aperture of the initial portion of the HL Tau flow causes the far side of the approaching lobe to be imaged at redshifted velocities close to the systemic. However, while this effect may explain the loss of the upper part of the large arc at blueshifted velocities, it would require the approaching lobe to have an opening angle larger than the receding one to explain the absence of a well-developed conical structure. A trace of faint conical arms of aperture angle similar to that in the red lobe is barely seen at the base of the blueshifted lobe, but it is not seen farther out, and it is covered by the signal of the large semicircular arc. This structure of the blue lobe may be due to HL Tau being located on the wall of a large-scale molecular bubble inflated by nearby XZ Tau, at 23″east of the source (Welch00; Yen19b). The expanding bubble may have compressed the outer envelope of HL Tau on its NE side, smashing the CO cavity, while the SW lobe remained unaffected. Signatures of such a large-scale bubble are visible at low blueshifted velocities in Fig. 2.
@@
In the following, we concentrate the analysis on the redshifted lobe, which presents a well-defined and relatively bright structure over a wide range of velocities. This allows a detailed inspection of the spatio-kinematic properties, for a meaningful comparison with proposed acceleration mechanisms. We defer the study of the fainter blueshifted lobe to a forthcoming investigation, adding that the squared 10″at the center of the region were discussed in Garufi22, where the kinematics of the rotating disk is analyzed from the same dataset, as well as possible accretion streamers.